In the 2004 summer olympics, the United states team won 3 more gold medals than the chinese team. If teh total number of gold medals won by both teams is 67, how many gold medals did each team win?
Answer by Mathtut(3670)

let the number of golds won by china and US be C and U respectively
:
C+U=67.......eq 1
U=C+3........eq 2
:
take U's value from eq 2 and plug it into eq 1
:
C+C+3=67
:
2C=64
:
C=32
:
U=C+3=32+3=35